OpenStudy (anonymous):

Medal and fan! Find the values of x and y that minimize the objective function C = 5x+8y for the constraints below. Just need someone to check my work :)

OpenStudy (anonymous):

Constraints: \[4x+3y \ge30\] \[x+3y\ge21\] \[x \ge0,y \ge0\]

OpenStudy (anonymous):

They tell me to write the equations in slope-intercept so I do: \[y \ge-\frac{ 4}{ 3 }x+10\] \[y \ge-\frac{ x }{ 3}+7\]

OpenStudy (anonymous):

Then they tell me to list the vertices so I use my graphing calculator: https://www.desmos.com/calculator/lu3gqjhp5v :)

OpenStudy (anonymous):

And if this is, is there a way to delete this? I really don't want any peeping toms who also need help on this to copy me >.<

OpenStudy (anonymous):

Your right

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Now to delete this ;)